For any cubic polynomial if 
f(1)=1
f(2)=2
f(3)=4
Then find 
f(6)
 

Grade:12th pass

1 Answers

Agrata Singh
5 years ago
Ket the cubic expression be x3+bx2+cx+d
f(1)=1+b+c+d=1
b+c+d=0..---(i)
 
f(2)=8+4b+2c+d=2
4b+2c+d+6=0..--(ii)
 
f(3)=27+9b+3c+d=4
9b+3c+d+23=0...--(iii)
 
Solve these eqns for b c d and put in to find f(x)
Hopefully this helps you